# UAR-NVC: A Unified AutoRegressive Framework for Memory-Efficient Neural Video Compression

### v1

@@ Abstract
Implicit Neural Representations (INRs) have demonstrated significant potential in video compression by representing videos as neural networks. However, as the number of frames increases, the memory consumption for training and inference increases substantially, posing challenges in resource-constrained scenarios. Inspired by the success of traditional video compression frameworks, which process video frame by frame and can efficiently compress long videos, we adopt this modeling strategy for INRs to decrease memory consumption, while aiming to unify the frameworks from the perspective of timeline-based autoregressive modeling. In this work, we present a novel understanding of INR models from an autoregressive (AR) perspective and introduce a Unified AutoRegressive Framework for memory-efficient Neural Video Compression (UAR-NVC). UAR-NVC integrates timeline-based and INR-based neural video compression under a unified autoregressive paradigm. It partitions videos into several clips and processes each clip using a different INR model instance, leveraging the advantages of both compression frameworks while allowing seamless adaptation to either in form. To further reduce temporal redundancy between clips, we treat the corresponding model parameters as proxies for these clips, and design two modules to optimize the initialization, training, and compression of these model parameters. In special, the Residual Quantization and Entropy Constraint (RQEC) module dynamically balances the reconstruction quality of the current clip and the newly introduced bitrate cost using the previously optimized parameters as conditioning. In addition, the Interpolation-based Initialization (II) module flexibly adjusts the degree of reference used during the initialization of neighboring video clips, based on their correlation. UAR-NVC supports adjustable latencies by varying the clip length. Extensive experimental results demonstrate that UAR-NVC, with its flexible video clip setting, can adapt to resource-constrained environments and significantly improve performance compared to different baseline models.
@@ II-A Traditional Video Compression Standards
Besides the H.26x series, the AVS series developed by the Audio and Video Coding Standard Workgroup of China [35] has also gained significant attention. Although the first version, AVS1, was much later than H.261, the latest version, AVS3, achieves performance comparable to H.266 in 2021. Meanwhile, the video coding standards VP8 [36], VP9 [37], and AV1 [38], led by Google, continue to evolve. Despite these advancements, nearly all of these standards follow the pipeline established by H.264, which underscores the robustness of the hybrid video coding framework. This highlights the stability and practicality of this framework, which served as an inspiration for our work.
@@ II-C INR for Image Compression
Subsequently, several works have focused on reducing the encoding time through meta-learning [18] and improved rate-distortion (RD) performance via better network architectures [51] or more precise quantization techniques [20, 52]. Other works have further enhanced the RD performance through model pruning[53] or by combining autoencoders [21]. COOL-CHIC [54] innovatively combines INR with feature embedding, proposing a low-complexity image codec. Building on COOL-CHIC, the subsequent work C3 [55] demonstrated RD performance close to VTM on the UVG video benchmark with less than 5k MACs/pixel for decoding. This highlights the promising advantages of implicit neural representation methods, particularly in resource-constrained environments.
@@ II-D INR for Video Compression
Different from coordinate-based INR models like SIREN [56] and COIN, NeRV [16] uses the time stamp $t$ as the input of the INR network, allowing it to reconstruct a entire video frame in one forward pass. This approach can achieve faster decoding speeds even than the practical traditional codec H.264. Following that, E-NeRV [22] combines spatial coordinates with time stamps and replaces the convolution kernel with two consecutive convolution kernels with smaller channels. PS-NeRV [57] reconstructs patches of the image instead of the whole image and uses Adaptive Instance Normalization (AdaIN) to modulate the features of the later convolution layers according to a time-coordinate condition. FF-NeRV [23] uses learnable multi-scale resolution feature grids to replace time or spatial coordinates, incorporates flow information into frame-wise representations to exploit the temporal redundancy across frames, and introduces quantization-aware training to shrink the quantization gap between training and model compression. HNeRV [24] trains an encoder to generate content-adaptive embeddings and proposes evenly distributed parameters for HNeRV, providing more capacity to store high-resolution content and video details. HiNeRV [25] uses the learnable feature grid not only at the network input but also in the subsequent upsampling module. By further improving the pruning and quantization parts of the model compression pipeline, HiNeRV achieves a $43.4\%$ overall bit rate saving over DCVC on the UVG dataset, measured in PSNR. Boost-NeRV [26] proposes three universal techniques for NeRV-style models in module conditioning, activation layers, and entropy constraints, respectively.
@@
Besides these excellent works that progressively improve the rate-distortion performance of Video INR, some works focus on the application of Video INR. To handle videos with frequent scene switching, D-NeRV [28] represents a large and diverse set of videos as a single neural network by using a multi-tiered structure together with a Visual Content Encoder and Motion-aware Decoder. Additionally, MVC [29] employs several spatial context enhancement and temporal correlation capturing techniques to further improve the representation capability of Video INR, achieving up to a $20\%$ bitrate reduction compared to the latest video coding standard H.266 in conference and surveillance videos. To process long and different resolution videos, NIRVANA [27] proposes fitting every continuous three frames using different independent patch-wise model instances and employs an autoregressive approach in model initialization and compression. However, NIRVANA is limited by its specific model structure, we unify the timeline-based framework and INR-based NVC framework from an autoregressive perspective. Our UAR-NVC framework allows users to choose different GOP sizes based on available resources, supports various INR base models, and includes two modules that significantly leverage the correlation between clips, thereby greatly improving rate-distortion (RD) performance for video compression.
@@ III-A Information Autoregressive
In the context where $x$ represents a pixel, PixelCNN [1] can be employed to model spatial relationships through the probability distribution $p(x_{i}|x_{i-1}x_{i-2}...)$, which utilizes preceding pixels to model the probability distribution of the current pixel. This approach is referred to as pixel-level AR. If $x$ denotes a frame, temporal relationships can be modeled using $p(x_{i}|x_{i-1})$ by leveraging motion estimation between neighboring frames [13]. This method captures temporal correlations within a video and is referred to as frame-level AR. Unlike generation tasks, compression tasks require recording both the signal distribution and the sampling position, as accurate signal recovery is necessary. Therefore, we consider the AR process in compression as an alternating process: initially obtaining the probability distribution, subsequently introducing new information decoded from the bitstream, and finally determining the value of the intermediate variable, which gradually converges to the original signal.
@@ III-B The Proposed UAR-NVC Framework
We begin by introducing the foundational concepts of our framework. A video with $T$ frames can be represented as $\mathbf{V}=\left\{\mathrm{v}_{t}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$. The INR model can be regarded as a mapping function $f_{\boldsymbol{\theta}}:\mathbb{R}\rightarrow\mathbb{R}^{H\times W\times 3}$ parameterized by $\boldsymbol{\theta}$. Given a time stamp $t$, the reconstructed frame $\hat{\mathrm{v}}_{t}$ is obtained, forming the complete reconstructed video $\hat{\mathbf{V}}=\left\{\hat{\mathrm{v}}_{t}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$. We define $p$ and $m$ as user-defined hyperparameters representing the size of GOP (Group of Pictures) and GOM (Group of Models), respectively. The video is partitioned based on these parameters. Each GOP instance is denoted as GOP#$\#k,k\in\mathbb{R}$, and each GOM instance as GOM#$\#k,k\in\mathbb{R}$, where $k$ represents the index of the respective GOP or GOM. The first model in a GOM is referred to as the I-model, meaning it does not depend on any preceding model. The remaining models in a GOM are termed P-models, as they reference the previous model for initialization and compression.
@@
Within a GOM, we employ an autoregressive approach for training and compressing each model. For Model#$\#0$, initialized with randomly parameters $\boldsymbol{\theta}_{0}$, the frames in GOP#$\#0$ serve as the training data. After training, the parameters are optimized from $\boldsymbol{\theta}_{0}$ to $\boldsymbol{\theta}_{0}^{*}$, denoted as: $\boldsymbol{\theta}_{0}\xrightarrow{train}\boldsymbol{\theta}_{0}^{*}$. The corresponding bitrate is defined as:
@@
Since the randomly initialized parameters $\boldsymbol{\theta}_{0}$ can be reconstructed using the initialization seed, and $\boldsymbol{\theta}_{0}^{*}$ can be read back from the bit stream, both (primarily $\boldsymbol{\theta}_{0}^{*}$) serve as reference information for processing Model#$\#1$. The relationship between adjacent GOPs is captured through model parameters. Specifically, instead of using randomly initialized parameters $\boldsymbol{\theta}_{1}$, we derive specially initialized parameters $\boldsymbol{\theta}_{1}^{\prime}$ for Model#$\#1$ by combining $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}_{0}^{*}$, defined as:
@@
This initialization strategy accelerates the training of Model#$\#1$, with experimental analyses provided in Sec. IV. Furthermore, since the final parameters $\boldsymbol{\theta}_{1}^{*}$ are obtained through training (or fine-tuning) from the primary parameters $\boldsymbol{\theta}_{1}^{\prime}$, denoted as: $\boldsymbol{\theta}_{1}^{\prime}\xrightarrow{train}\boldsymbol{\theta}_{1}^{*}$, we store the difference between $\boldsymbol{\theta}_{1}^{\prime}$ and $\boldsymbol{\theta}_{1}^{*}$ instead of directly recording $\boldsymbol{\theta}_{1}^{*}$. The bitrate of Model#$\#1$ is given by:
@@ III-C Group of Models
Traditional video compression standards, such as H.264, categorize video frames into three types: intra-coded (I) frames, predictive-coded (P) frames, and bidirectionally predicted (B) frames. One purpose of this classification is to regulate the reference direction of each frame. The Group of Pictures (GOP) structure segments a video sequence into multiple groups, where frames within a GOP can reference each other while ensuring a one-way reference direction between any two frames. Across different GOPs, compression and decompression occur independently, enabling resynchronization during decoding in practical applications such as video jump playback.
@@ III-D Residual Quantization and Entropy Constrain
$\begin{cases}Loss_{d}=MSE\left(\mathrm{v},f\left(t;\boldsymbol{\theta}_{k}^{*}\right)\right),\\ Loss_{r}=I\left(\boldsymbol{\theta}_{k}^{*}-\boldsymbol{\theta}_{k}^{\prime}\right),\\ Loss=Loss_{r}+\lambda\cdot Loss_{d},\\ \end{cases}$ (7)
@@ III-E Interpolation-based Model Initialization
Since adjacent frames in a video exhibit strong correlations, autoregressively initialized models that leverage parameters from neighboring model perform well when $p$ is small. However, as $p$ increases, these correlations weaken or even disappear. Consequently, the direct initialization (D-init) method may perform poorly, potentially even worse than random initialization (R-init). This phenomenon may arise from the large space of network parameters. The optimal parameters for GOP#$\#a$ may be significantly different from those for GOP#$\#b$, both of which are also distant from the randomly initialized parameters. In extreme cases, where the optimal parameters of two models lie in opposite directions, D-init can increase optimization difficulty rather than providing a useful starting point.
@@
To determine the optimal $\varepsilon$, we analyze the rate-distortion (RD) performance across different values of $\varepsilon$ and $p$. Our findings indicate that the best $\varepsilon$ varies depending on $p$. Specifically, as $p$ increases, the gap between adjacent GOPs widens, necessitating a larger $\varepsilon$ to achieve the best RD performance. This trend likely arises because a larger $p$ results in greater differences between corresponding frames in successive GOPs. To quantify the gap between two GOPs, we measure the Mean Squared Error (MSE) of corresponding frames [27]. To establish a numerical relationship between MSE and $\varepsilon$, we conduct a statistical analysis of the optimal $\varepsilon$ and MSE values across different $p$ settings using three sequences from the MCL-JCV [59] dataset, which are non-repetitive with the test datasets we use in Sec. IV. We assume this function follows a variant of an exponential function:
@@ IV-A1 Test Sequences and Base Models
We select HNeRV [24] and HiNeRV [25] as the base models for our framework and evaluate them on the UVG dataset and VVC Class B dataset. For the UVG dataset, we use six sequences (“HoneyBee”, “Bosphorus”, “Beauty”, “YachtRide”, “Jockey”, “ReadySetGo”), each containing 600 frames. We set $m=5$ and evaluate $p$ with values of (6, 30, 120) for the UVG dataset and (5, 25, 100) for the VVC Class B dataset to evaluate small, medium, and large $p$ scenarios. We provide two versions of the base models. The “base” version employs the QAT [25] and D init. The “ours” version incorporates our RQEC and II modules.

### v2

@@ Abstract
Implicit Neural Representations (INRs) have demonstrated significant potential in video compression by representing videos as neural networks. However, as the number of frames increases, the memory consumption for training and inference increases substantially, posing challenges in resource-constrained scenarios. Inspired by the success of traditional video compression frameworks, which process video frame by frame and can efficiently compress long videos, we adopt this modeling strategy for INRs to decrease memory consumption, while aiming to unify the frameworks from the perspective of timeline-based autoregressive modeling. In this work, we present a novel understanding of INR models from an autoregressive (AR) perspective and introduce a Unified AutoRegressive Framework for memory-efficient Neural Video Compression (UAR-NVC). UAR-NVC integrates timeline-based and INR-based neural video compression under a unified autoregressive paradigm. It partitions videos into several clips and processes each clip using a different INR model instance, leveraging the advantages of both compression frameworks while allowing seamless adaptation to either in form. To further reduce temporal redundancy between clips, we treat the corresponding model parameters as proxies for these clips, and design two modules to optimize the initialization, training, and compression of these model parameters. In special, the Residual Quantization and Entropy Constraint (RQEC) module dynamically balances the reconstruction quality of the current clip and the newly introduced bitrate cost using the previously optimized parameters as conditioning. In addition, the Interpolation-based Initialization (II) module flexibly adjusts the degree of reference used during the initialization of neighboring video clips, based on their correlation. UAR-NVC supports adjustable latencies by varying the clip length. Extensive experimental results demonstrate that UAR-NVC, with its flexible video clip setting, can adapt to resource-constrained environments and significantly improve performance compared to different baseline models. The project page: https://wj-inf.github.io/UAR-NVC-page/.
@@ II-A Traditional Video Compression Standards
Besides the H.26x series, the AVS series developed by the Audio and Video Coding Standard Workgroup of China [35] has also gained significant attention. Meanwhile, the video coding standards VP8 [36], VP9 [37], and AV1 [38], led by Google, continue to evolve. Despite these advancements, nearly all of these standards follow the pipeline established by H.264, which underscores the robustness of the hybrid video coding framework. This highlights the stability and practicality of this framework, which served as an inspiration for our work.
@@ II-C INR for Image Compression
Subsequently, several works have focused on reducing the encoding time through meta-learning [18] and improved rate-distortion (RD) performance via better network architectures [51] or more precise quantization techniques [20, 52]. Other works have further enhanced the RD performance through model pruning[53] or by combining autoencoders [21]. COOL-CHIC [54] innovatively combines INR with feature embedding, proposing a low-complexity image codec. Building on COOL-CHIC, the subsequent work C3 [55] further optimized its quantization strategies and model architecture, demonstrating RD performance close to VTM on the UVG video benchmark with less than 5k MACs/pixel for decoding. Besides, C3 models the video as 3D patches to mitigate the large memory footprint associated with full-video processing. This highlights the promising advantages of implicit neural representation methods, particularly in resource-constrained environments.
@@ II-D INR for Video Compression
Different from coordinate-based INR models like SIREN [56] and COIN, NeRV [16] uses the time stamp $t$ as the input of the INR network, allowing it to reconstruct a entire video frame in one forward pass. This approach can achieve faster decoding speeds even than the practical traditional codec H.264. IPF [57] using the explicit relation of two frames (I-frame and P-frame) to facilitate modeling. E-NeRV [22] combines spatial coordinates with time stamps and replaces the convolution kernel with two consecutive convolution kernels with smaller channels. PS-NeRV [58] reconstructs patches of the image instead of the whole image and uses Adaptive Instance Normalization (AdaIN) to modulate the features of the later convolution layers according to a time-coordinate condition. FF-NeRV [23] uses learnable multi-scale resolution feature grids to replace time or spatial coordinates, incorporates flow information into frame-wise representations to exploit the temporal redundancy across frames, and introduces quantization-aware training to shrink the quantization gap between training and model compression. HNeRV [24] trains an encoder to generate content-adaptive embeddings and proposes evenly distributed parameters for HNeRV, providing more capacity to store high-resolution content and video details. DSNeRV [59] decomposes video into static part and dynamic part, using two group embedding to represent each. T-NeRV [60] decomposes embedding into frame level and GOP level, and use the encoder in HNeRV to learning better frame embedding. HiNeRV [25] uses the learnable feature grid not only at the network input but also in the subsequent upsampling module. By further improving the pruning and quantization parts of the model compression pipeline, HiNeRV achieves a $43.4\%$ overall bit rate saving over DCVC on the UVG dataset, measured in PSNR. Boost-NeRV [26] proposes three universal techniques for NeRV-style models in module conditioning, activation layers, and entropy constraints, respectively.
@@
Besides these excellent works that progressively improve the rate-distortion performance of Video INR, some works focus on the application of Video INR. To handle videos with frequent scene switching, D-NeRV [28] represents a large and diverse set of videos as a single neural network by using a multi-tiered structure together with a Visual Content Encoder and Motion-aware Decoder. Additionally, MVC [29] employs several spatial context enhancement and temporal correlation capturing techniques to further improve the representation capability of Video INR, achieving up to a $20\%$ bitrate reduction compared to the latest video coding standard H.266 in conference and surveillance videos. PNVC [61] introduces a pre-trained image encoder to accelerate the encoding process. To process long and different resolution videos, NIRVANA [27] proposes fitting every continuous three frames using different independent patch-wise model instances and employs an autoregressive approach in model initialization and compression. However, NIRVANA is limited by its specific model structure, we unify the timeline-based framework and INR-based NVC framework from an autoregressive perspective. Our UAR-NVC framework allows users to choose different GOP sizes based on available resources, supports various INR base models, and includes two modules that significantly leverage the correlation between clips, thereby greatly improving rate-distortion (RD) performance for video compression.
@@ III-A Information Autoregressive
[…]
@@ III-B The Proposed UAR-NVC Framework
[…]
@@
Within a GOM, we employ an autoregressive approach for training and compressing each model. For Model#0\#0, initialized with randomly parameters $\boldsymbol{\theta}_{0}$, the frames in GOP#0\#0 serve as the training data. After training, the parameters are optimized from $\boldsymbol{\theta}_{0}$ to $\boldsymbol{\theta}_{0}^{*}$, denoted as: $\boldsymbol{\theta}_{0}\xrightarrow{train}\boldsymbol{\theta}_{0}^{*}$. The corresponding bitrate is defined as:
@@
Since the randomly initialized parameters $\boldsymbol{\theta}_{0}$ can be reconstructed using the initialization seed, and $\boldsymbol{\theta}_{0}^{*}$ can be read back from the bit stream, both (primarily $\boldsymbol{\theta}_{0}^{*}$) serve as reference information for processing Model#1\#1. The relationship between adjacent GOPs is captured through model parameters. Specifically, instead of using randomly initialized parameters $\boldsymbol{\theta}_{1}$, we derive specially initialized parameters $\boldsymbol{\theta}_{1}^{\prime}$ for Model#1\#1 by combining $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}_{0}^{*}$, defined as:
@@
This initialization strategy accelerates the training of Model#1\#1, with experimental analyses provided in Sec. IV. Furthermore, since the final parameters $\boldsymbol{\theta}_{1}^{*}$ are obtained through training (or fine-tuning) from the primary parameters $\boldsymbol{\theta}_{1}^{\prime}$, denoted as: $\boldsymbol{\theta}_{1}^{\prime}\xrightarrow{train}\boldsymbol{\theta}_{1}^{*}$, we store the difference between $\boldsymbol{\theta}_{1}^{\prime}$ and $\boldsymbol{\theta}_{1}^{*}$ instead of directly recording $\boldsymbol{\theta}_{1}^{*}$. The bitrate of Model#1\#1 is given by:
@@ III-C Group of Models
Traditional video compression standards, such as H.264, categorize video frames into three types: intra-coded (I) frames, predictive-coded (P) frames, and bidirectionally predicted (B) frames, to regulate the reference direction of each frame. The Group of Pictures (GOP) structure segments a video sequence into multiple groups, where frames within a GOP can reference each other while ensuring a one-way reference direction between any two frames. Across different GOPs, compression and decompression occur independently, enabling resynchronization during decoding in practical applications such as video jump playback.
@@ III-D Residual Quantization and Entropy Constrain
[…]
@@
We design this module in the principle of local temporal similarity within video sequences. It utilizes the model parameters of the preceding GOP as a reference for the current one. To enhance RD performance, our RQEC module establish a more accurate rate estimation model through consistent entropy estimation.
@@ III-E Interpolation-based Model Initialization
Since adjacent frames in a video exhibit strong correlations, autoregressively initialized models that leverage parameters from neighboring model perform well when $p$ is small. However, as $p$ increases, these correlations weaken or even disappear. Consequently, the direct initialization (D-init) method may perform poorly, potentially even worse than random initialization (R-init). This phenomenon may arise from the large space of network parameters. The optimal parameters for GOP#a\#a may be significantly different from those for GOP#b\#b, both of which are also distant from the randomly initialized parameters. In extreme cases, where the optimal parameters of two models lie in opposite directions, D-init can increase optimization difficulty rather than providing a useful starting point.
@@
In Fig. 5, we test the RD performance of HiNeRV on “ReadySetGo” with different random percent, which indicates that the optimal $\varepsilon$ value corresponding to different gops is different. To determine the optimal $\varepsilon$, we analyze the rate-distortion (RD) performance across different values of $\varepsilon$ and $p$. Our findings indicate that the best $\varepsilon$ varies depending on $p$. Specifically, as $p$ increases, the gap between adjacent GOPs widens, necessitating a larger $\varepsilon$ to achieve the best RD performance. This trend likely arises because a larger $p$ results in greater differences between corresponding frames in successive GOPs. To quantify the gap between two GOPs, we measure the Mean Squared Error (MSE) of corresponding frames [27]. To establish a numerical relationship between MSE and $\varepsilon$, we conduct a statistical analysis of the optimal $\varepsilon$ and MSE values across different $p$ settings using three sequences from the MCL-JCV [63] dataset, which are non-repetitive with the test datasets we use in Sec. IV. We assume this function follows a variant of an exponential function:
@@
The adaptive can be viewed as a simple, “implicitly learned” meta-strategy, embodying the idea of “learning how to initialize.” It dynamically decides the extent to which past experience ($\boldsymbol{\theta}_{k-1}^{*}$) should be trusted and utilized, based on the similarity (measured by MSE) between the current task (encoding GOP k) and the previous task (encoding GOP k-1).
@@ IV EXPERIMENTS
TABLE I: BD-rate (%) for PSNR/MS-SSIM on different datasets and $p$ (GOP).
Method Anchor UVG VVC Class B
PSNR MS-SSIM PSNR MS-SSIM
GOP6 GOP30 GOP120 GOP6 GOP30 GOP120 GOP5 GOP25 GOP100 GOP5 GOP25 GOP100
NeRV (ours) NeRV (base) -45.04 -29.64 -35.99 -39.73 -24.15 -30.59 -34.73 -30.99 -36.84 -31.92 -27.02 -32.29
HNeRV (ours) HNeRV (base) -41.08 -31.89 -11.79 -19.42 -30.26 -19.03 -27.85 -36.18 -32.90 -22.78 -34.44 -28.64
HiNeRV (ours) HiNeRV (base) -58.94 -61.57 -64.97 -51.27 -53.74 -56.12 -49.81 -52.24 -58.12 -44.11 -42.20 -53.54
Avg. -48.35 -41.03 -37.58 -36.81 -36.05 -35.25 -37.46 -39.80 -42.62 -32.94 -34.55 -38.16
HiNeRV (ours) X265 (veryslow) 20.43 -17.60 -34.29 -30.86 -51.29 -53.79 85.55 36.34 -18.79 3.01 -4.82 -50.57
@@ IV-A1 Test Sequences and Base Models
To validate the effectiveness of our framework, we conducted comprehensive evaluations on multiple models, including NeRV [16], HNeRV [24] and HiNeRV [25]. These tests spanned diverse datasets characterized by varied video content, resolution, and length:
  • UVG dataset [64]: Six sequences (“HoneyBee”, “Bosphorus”, “Beauty”, “YachtRide”, “Jockey”, “ReadySetGo”), each comprising 600 frames at 1920x1080 resolution.
  • VVC Class B dataset: Five sequences (“BasketballDrive”, “BQTerrace”, “Cactus”, “MarketPlace”, “RitualDance”), each containing 500 frames at 1920x1024 resolution.
  • Big Buck Bunny: A video sequence of 132 frames with a 1280x720 resolution, used in NeRV.
We set $m=5$ and evaluate $p$ with values of (6, 30, 120) for the UVG dataset and (5, 25, 100) for the VVC Class B dataset to evaluate small, medium, and large $p$ scenarios. For bunny dataset, we set $m=22$ and $p\in(6,22)$. We provide two versions of the base models. The “base” version employs the QAT [25] and D init. The “ours” version incorporates our RQEC and II modules.